# Via P. Giuria 1, 10125 Torino, Italy

### v1

@@ 1 Introduction and discussion
The central idea of the conformal bootstrap program is to constrain the dynamics of conformal field theories (CFTs) using symmetries and internal consistency. After the breakthrough of Rattazzi:2008pe , this program was revived, leading to important numerical and analytical results (see Rychkov:2023wsd ; Bissi:2022mrs and references therein). One of the early achievements of the analytic bootstrap is due to Komargodski:2012ek ; Fitzpatrick:2012yx , where it was found that a generic CFT must contain an infinite tower of double-twist operators, whose dimension approaches the generalized free field (GFF) dimension at large spin. The main idea leading to this result was to exploit the crossing equation for the four-point function of identical operators in Lorentzian signature and consider its behaviour in the lightcone limit, i.e. when two operators become lightlike separated. Several important developments followed, either trying to make the statement mathematically rigorous Qiao:2017xif ; Kravchuk:2020scc ; Kravchuk:2021kwe ; Pal:2022vqc ; vanRees:2024xkb or extending and exploiting the initial result Fitzpatrick:2014vua ; Alday:2015eya ; Kaviraj:2015cxa ; Li:2015itl ; Komargodski:2016gci ; Alday:2016njk ; Simmons-Duffin:2016wlq ; Pal:2023cgk . In particular, the Lorentzian inversion formula for four-point correlators encoded the initial observation of Komargodski:2012ek ; Fitzpatrick:2012yx in a general framework which allows to extract the CFT data from the discontinuity of the correlator, thus providing an efficient computational method for all those theories that admit a perturbative expansion and some control on the perturbative spectrum Caron-Huot:2017vep .
@@
In this paper, we analyze a setup that enlarges the set of available CFT data with a milder complication of the kinematical structure. To achieve this, we consider a CFT in the presence of a conformal defect, i.e. an extended excitation which preserves conformal invariance along its profile. Starting with the seminal work of Billo:2016cpy , the application of bootstrap techniques to conformal defects has evolved into an independent research line, the defect bootstrap program. If one is interested in the interaction between bulk and defect degrees of freedom, the first observable with a non-trivial crossing equation is the two-point function of bulk operators. The lack of positivity in the bulk channel does not allow for a straightforward application of numerical bootstrap methods (see however Meineri:2025qwz ; Lanzetta:2025xfw for recent progress), but it is certainly possible to apply analytic bootstrap techniques. In particular, Lorentzian inversion formulae Lemos:2017vnx ; Liendo:2019jpu and dispersion relations Bianchi:2022ppi ; Barrat:2022psm have been derived for both OPE channels and their simplicity allowed for the computation of the correlator in situations where one can expand in a small paremeter having some control on the spectrum of exchanged operators Bissi:2018mcq ; Dey:2020jlc ; Barrat:2021yvp ; Liendo:2016ymz ; Gimenez-Grau:2019hez ; Gimenez-Grau:2021wiv ; Bianchi:2022sbz ; Gimenez-Grau:2022ebb ; Bianchi:2023gkk ; Chen:2023yvw ; Chen:2024orp .
@@
Here we consider the correlator $\big\langle\Phi\Phi\hat{\phi}\big\rangle$, where we added a defect operator $\hat{\phi}$. This correlator admits two different OPE expansions, which are schematically represented in Figure 2. We study the resulting crossing equation in the limit where the two bulk operators become lightlike separated as shown in Figure 2 through a red curly line. In the bulk channel, the exchanged operator with lowest twist provides the leading contribution in the lightcone limit. A crucial difference compared to the BB case is that the bulk identity operator is not exchanged as it does not couple with a generic defect operator $\hat{\phi}$ (unless $\hat{\phi}$ is the defect identity of course). Therefore, to understand which operators dominate the bulk expansion at low twist we need some insight into the twist spectrum of the bulk theory. Focusing on unitary theories, we know that the unitarity bounds for the bulk theory impose $h\geq\frac{d-2}{4}$ for scalars and $h\geq\frac{d-2}{2}$ for spinning operators (remember that $h=\frac{\Delta-J}{2}$ is the half twist, so for the stress tensor $h=\frac{d-2}{2}$ and for a scalar $h=\frac{\Delta}{2}$). In this work we exclude higher-spin conserved currents as they are expected to be absent in an interacting CFT for $d>2$ Maldacena:2011jn . Therefore, there are no spinning operators saturating the unitarity bound for $J>2$. The leading-twist contribution is given either by the stress tensor operator or by a light scalar operator (with $\frac{d-2}{2}\leq\Delta\leq d-2$) [^2] Strictly speaking, one should consider also a spin-one conserved current, but the two-point function of a bulk current with a defect scalar operator vanishes for generic codimension. It could be non-vanishing for specific cases where parity-odd contributions must be included, see e.g. Bianchi:2019sxz . In this work we analyze both these options and we study the implications for the defect spectrum at large spin.
@@
Requiring consistency of the crossing equation in the lightcone limit we can find an expression for the large-spin beahvior of the bulk-to-defect couplings $b_{\Phi[\hat{\phi}\Phi]}$ and $b_{\Phi[\hat{\phi}\hat{\phi}\Phi]}$ in terms of the bulk OPE data, i.e. the product of a bulk OPE coefficient $\lambda_{\Phi\Phi\mathcal{O}_{\star}}$ and a bulk-to-defect coupling $b_{\mathcal{O}_{\star}\hat{\phi}}$, where $\mathcal{O}_{\star}$ is the leading-twist exchanged operator, i.e. the stress tensor or a light scalar operator. Finding closed-form expressions for these couplings at arbitrary $m_{1}$ and $m_{2}$ is not easy, but we provide a relation that allows to extract these OPE coefficients recursively in $m_{1}$ and $m_{2}$. Furthermore, thanks to the simplicity of the defect setup, we can obtain closed-form expressions for the cases $m_{1}=0$ or $m_{2}=0$ and arbitrary values of the other index. Here we show only the result for a scalar operator. More general expressions for arbitrary spin can be found in Section 3. For a scalar bulk exchange of dimension $\Delta=2h_{\star}$ we find that the leading behaviour for the defect CFT data at large spin is
@@ A BBD correlator can be expressed as follows
$\mathcal{G}(z,\bar{z},x)=\sum_{\Delta,J,n}\lambda_{\Phi\Phi\mathcal{O}}\,b_{\mathcal{O}\hat{\phi}}^{(n)}\,f_{\Delta,J,n}\,,(z,\bar{z},x)$ (26)
@@ Acknowledgements.
The authors would like to thank M.Lemos, M. Meineri and V. Schomerus for useful discussions. LB’s and LQ’s research is partially supported by the MUR PRIN contract 2022N9CTAE "Constraining strongly coupled quantum field theories using symmetry". LB and AM would like to thank the Isaac Newton Institute for Mathematical Sciences, Cambridge, for support and hospitality during the programme "Quantum Field Theory with Boundaries, Impurities, and Defects" where work on this paper was undertaken. This work was supported by EPSRC grant no EP/R014604/1.
@@ Appendix D Large spin expansion of defect-channel blocks
Combining this result with (36) and (104), performing the sum with index $k$, and re-expressing variables in terms of the cross-ratios $(x,z,\bar{z})$, we obtain the leading behavior of the of the defect-channel conformal blocks in the simultaneous lightcone/large-spin limit:

### v2

@@ 1 Introduction and discussion
The central idea of the conformal bootstrap program is to constrain the dynamics of conformal field theories (CFTs) using symmetries and internal consistency. After the breakthrough of Rattazzi:2008pe , this program was revived, leading to important numerical and analytical results (see Rychkov:2023wsd ; Bissi:2022mrs and references therein). One of the early achievements of the analytic bootstrap is due to Komargodski:2012ek ; Fitzpatrick:2012yx , where it was found that a generic CFT must contain an infinite tower of double-twist operators, whose dimension approaches the generalized free field (GFF) dimension at large spin. The main idea leading to this result was to exploit the crossing equation for the four-point function of identical operators in Lorentzian signature and consider its behaviour in the lightcone limit, i.e. when two operators become lightlike separated. Several important developments followed, either trying to make the statement mathematically rigorous Qiao:2017xif ; Kravchuk:2020scc ; Kravchuk:2021kwe ; Pal:2022vqc ; vanRees:2024xkb or extending and exploiting the initial result Fitzpatrick:2014vua ; Alday:2015eya ; Kaviraj:2015cxa ; Li:2015itl ; Komargodski:2016gci ; Alday:2016njk ; Simmons-Duffin:2016wlq ; Kusuki:2021gpt ; Numasawa:2022cni ; Pal:2023cgk . In particular, the Lorentzian inversion formula for four-point correlators encoded the initial observation of Komargodski:2012ek ; Fitzpatrick:2012yx in a general framework which allows to extract the CFT data from the discontinuity of the correlator, thus providing an efficient computational method for all those theories that admit a perturbative expansion and some control on the perturbative spectrum Caron-Huot:2017vep .
@@
In this paper, we analyze a setup that enlarges the set of available CFT data with a milder complication of the kinematical structure. To achieve this, we consider a CFT in the presence of a conformal defect, i.e. an extended excitation which preserves conformal invariance along its profile. Starting with the seminal work of Billo:2016cpy , the application of bootstrap techniques to conformal defects has evolved into an independent research line, the defect bootstrap program. If one is interested in the interaction between bulk and defect degrees of freedom, the first observable with a non-trivial crossing equation is the two-point function of bulk operators. The lack of positivity in the bulk channel does not allow for a straightforward application of numerical bootstrap methods (see however Meineri:2025qwz ; Lanzetta:2025xfw for recent progress), but it is certainly possible to apply analytic bootstrap techniques. In particular, Lorentzian inversion formulae Lemos:2017vnx ; Liendo:2019jpu and dispersion relations Bianchi:2022ppi ; Barrat:2022psm have been derived for both OPE channels and their simplicity allowed for the computation of the correlator in situations where one can expand in a small parameter having some control on the spectrum of exchanged operators Bissi:2018mcq ; Dey:2020jlc ; Barrat:2021yvp ; Liendo:2016ymz ; Gimenez-Grau:2019hez ; Gimenez-Grau:2021wiv ; Bianchi:2022sbz ; Gimenez-Grau:2022ebb ; Bianchi:2023gkk ; Chen:2023yvw ; Chen:2024orp .
@@
Here we consider the correlator $\big\langle\Phi\Phi\hat{\phi}\big\rangle$, where we added a defect operator $\hat{\phi}$. This correlator admits two different OPE expansions, which are schematically represented in Figure 2. We study the resulting crossing equation in the limit where the two bulk operators become lightlike separated as shown in Figure 2 through a red curly line. In the bulk channel, the exchanged operator with lowest twist provides the leading contribution in the lightcone limit. A crucial difference compared to the BB case is that the bulk identity operator is not exchanged as it does not couple with a generic defect operator $\hat{\phi}$ (unless $\hat{\phi}$ is the defect identity of course). Therefore, to understand which operators dominate the bulk expansion at low twist we need some insight into the twist spectrum of the bulk theory. Focusing on unitary theories, we know that the unitarity bounds for the bulk theory impose $h\geq\frac{d-2}{4}$ for scalars and $h\geq\frac{d-2}{2}$ for spinning operators (remember that $h=\frac{\Delta-J}{2}$ is the half twist, so for the stress tensor $h=\frac{d-2}{2}$ and for a scalar $h=\frac{\Delta}{2}$). In this work we exclude higher-spin conserved currents as they are expected to be absent in an interacting CFT for $d>2$ Maldacena:2011jn . Therefore, there are no spinning operators saturating the unitarity bound for $J>2$. The leading-twist contribution is given either by the stress tensor operator or by a light scalar operator (with $\frac{d-2}{2}\leq\Delta\leq d-2$) [^2] Strictly speaking, one should consider also a spin-one conserved current, but the two-point function of a bulk current with a defect scalar operator vanishes for generic codimension. It could be non-vanishing for specific cases where parity-odd contributions must be included, see e.g. Bianchi:2019sxz ; Bachas:2025brp . In this work we analyze both these options and we study the implications for the defect spectrum at large spin.
@@
Requiring consistency of the crossing equation in the lightcone limit we can find an expression for the large-spin behavior of the bulk-to-defect couplings $b_{\Phi[\hat{\phi}\Phi]}$ and $b_{\Phi[\hat{\phi}\hat{\phi}\Phi]}$ in terms of the bulk OPE data, i.e. the product of a bulk OPE coefficient $\lambda_{\Phi\Phi\mathcal{O}_{\star}}$ and a bulk-to-defect coupling $b_{\mathcal{O}_{\star}\hat{\phi}}$, where $\mathcal{O}_{\star}$ is the leading-twist exchanged operator, i.e. the stress tensor or a light scalar operator. Finding closed-form expressions for these couplings at arbitrary $m_{1}$ and $m_{2}$ is not easy, but we provide a relation that allows to extract these OPE coefficients recursively in $m_{1}$ and $m_{2}$. Furthermore, thanks to the simplicity of the defect setup, we can obtain closed-form expressions for the cases $m_{1}=0$ or $m_{2}=0$ and arbitrary values of the other index. Here we show only the result for a scalar operator. More general expressions for arbitrary spin can be found in Section 3. For a scalar bulk exchange of dimension $\Delta=2h_{\star}$ we find that the leading behavior for the defect CFT data at large spin is
@@ A BBD correlator can be expressed as follows
$\mathcal{G}(z,\bar{z},x)=\sum_{\Delta,J,n}\lambda_{\Phi\Phi\mathcal{O}}\,b_{\mathcal{O}\hat{\phi}}^{(n)}\,f_{\Delta,J,n}\,(z,\bar{z},x)$ (26)
@@ Acknowledgements.
The authors would like to thank M. Lemos, M. Meineri and V. Schomerus for useful discussions. LB’s and LQ’s research is partially supported by the MUR PRIN contract 2022N9CTAE "Constraining strongly coupled quantum field theories using symmetry". LB and AM would like to thank the Isaac Newton Institute for Mathematical Sciences, Cambridge, for support and hospitality during the programme "Quantum Field Theory with Boundaries, Impurities, and Defects" where work on this paper was undertaken. This work was supported by EPSRC grant no EP/R014604/1. The authors participate in the project HORIZON-MSCA-2023-SE-01-101182937HeI.
@@ Appendix D Large spin expansion of defect-channel blocks
Combining this result with (36) and (104), performing the sum with index $k$, and re-expressing variables in terms of the cross-ratios $(x,z,\bar{z})$, we obtain the leading behavior of the defect-channel conformal blocks in the simultaneous lightcone/large-spin limit: